I do not understand these type of questions!...
one number is 6 less than another. Three times the smaller is 2 more than twice the larger. Find the numbers.
Bill has twice as much money as Bob. Paul has $12 more than Bill. Togethere they have $92. How much money does Bob have?
Find two consectutive numbers that total 94.

Answer by stanbon(48558)

One number is 6 less than another. Three times the smaller is 2 more than twice the larger. Find the numbers.
Let the smaller number be y; the larger be x.
EQUATION:
1st: x-y=6
2nd: 3y-2=2x
Solving by substitution solve 1st by x:
3rd: x=y+6
Substitute into 2nd to get:
3y-2=2(y+6)
3y-2=2y+12
y=14
x=y+6=20
SOLUTION: x=20; y=14
--------------------------------------------------
Bill has twice as much money as Bob. Paul has $12 more than Bill. Togethere they have $92. How much money does Bob have?
Let amount Bob has be "x" dollars.
Then Bill has "2x" dollars.
And Paul has "2x+12" dollars.
EQUATION:
x+2x+2x+12=92 dollars
4x=80
x=20 Dollars (amount Bob has)
2x=40 Dollars (amount Bill has)
2x+12 Dollars (amount Paul has)
-------------------------------
Find two consecutive numbers that total 94.
Let one number be "x"
The next consecutive number is "x+1"
EQUATION:
x+x+1=94
2x+1=94
2x=93
x=46.5
x+1=47.5
These numbers do not satisfy the normal definition
of "consecutive numbers". Consecutive numbers implies
that one is odd, the other is even. An odd number
added to an even number is always odd, so the sum
could not be 94.
Cheers,
Stan H.
Stan H.